Felix and Freddie are married with annual taxable income of $230,000. They pay income tax according to the following schedule: Over--------But Not Over---------Tax Rate $0.................$43,850...................15% $43,850..........$105,950..................??? $105,950........ $361,450.................31% If the total personal income tax they pay is $59,316, which of the following comes closest to the tax rate for income between $43,850 and 105,950 (the middle tax rate)?.

a.

23%

b.

25%

c.

24%

d.

21%

e.

22%
Answer #1
Tax paid for first $43850 6577.5 =43850*15%
Taxable income between $105,950-$361,450 124050 =230000-105950
Taxable payable between $105,950-$361,450 38455.5 =124050*31%
Remaining tax payable 14283 =59316-6577.5-38455.5
Divide by Remaining taxable income 62100 =105950-43850
Tax rate between $43,850 and 105,950 23%
Option A 23% is correct